40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jan 9 09:18:47 2017, Steve, Update, PEM, power glitch MondayMorning.png
    Reply  Tue Jan 10 16:20:11 2017, Steve, Update, CDS, power glitch......Raid is rebuilding rebuilding_in_progress.png
    Reply  Tue Jan 10 21:47:00 2017, rana, Update, CDS, power glitch 
       Reply  Tue Jan 10 22:55:43 2017, gautam, Update, CDS, power glitch - recovery steps 
          Reply  Wed Jan 11 16:35:03 2017, gautam, Update, CDS, power glitch - recovery progress powerGlitchRecoveryJan2017.png
             Reply  Thu Jan 12 17:31:51 2017, gautam, Update, CDS, DC errors DCerror.png
                Reply  Fri Jan 13 21:41:23 2017, Koji, Update, CDS, DC errors 
Message ID: 12708     Entry time: Thu Jan 12 17:31:51 2017     In reply to: 12702     Reply to this: 12715
Author: gautam 
Type: Update 
Category: CDS 
Subject: DC errors 

The IFO is more or less back to an operational state. Some details:

  1. The IMC mirror excess motion alluded to in the previous elog was due to some timing issues on c1sus. The "DAC" and "DK" blocks in the c1x02 diag word were red instead of green. Restarting all the models on c1sus fixed the problem
  2. When c1ioo was restarted, all of Koji's changes (digital) to the MC WFS servo where lost as they were not committed to the SDF. Eric suggested that I could just restore them from burt snapshots, which is what I did. I used the c1iooepics.snap file from 12:19PM PST on 26 December 2016, which was a time when the WFS servo was working well as per this elog by Koji. I have also committed all the changes to the SDF. IMC alignment has been stable for the last 4 hours.
  3. Johannes aligned and locked the arms today. There was a large DC offset on POX11, which was zeroed out by closing the PSL shutter and running LSC offsets. Both arms lock and stay aligned now.
  4. The doubling oven controller at the Y end was switched off. Johannes turned it on.
  5. Eric and I started a data consistency check on the RAID array yesterday, it has completed today and indicated no issues
  6. NDS2 is now running again on megatron so channel access from outside should(???) be possible again.

One error persists - the "DC" indicator (data concentrator?) on the CDS medm screen for the various models spontaneously go red and return to green often. Is this a known issue with an easy fix?

ELOG V3.1.3-